OpenStudy (unofficialllyy):

Imagine 200kg a rocket being launched. While in the air, it is experiencing a 5,000N thrust force as well as a 1,000N drag force. What is the acceleration of the rocket?

OpenStudy (ljetibo):

$$F = m*a$$ when dealing with several forces: $$F_1+F_2+F_3\dots + F_n = m*a$$ or in a shorter notation: $$\sum_0^n F_i = m*a$$ easy peasy

OpenStudy (osprey):

A spot of rocket science, yum yum. 5000N thrust is force from fuel to move the rocket. 1000N of drag is a resisting force caused by friction between the air and the rocket. Resultant force, is 4000N of "useable" thrust. (1000N to overcome the air resistance drag, 4000N to accelerate rocket). Mass of rocket is 200kg. The mathematical form of Newton's second law of motion kicks in here - F = ma. 4000 = 200xa. So, the acceleration is 4000/200 =20 metres per second each second. (If my arithmetic has worked, that is). Acceleration of gravity is 10 metres per second each second, so the rocket, plus contents, is being subjected to "2g", as they say at NASA and other places in USA.. This thrust comes from the rocket fuel. So, the longer the massive thing does its trick of accelerating at 2g, it's burning rocket fuel. And that means that its mass is decreasing. May come a point in the journey when it's burned all the fuel from one of the fuel tanks ("stages" in NASA speak ?). So, it's jettisoned. I don't know if the numbers quoted in the question represent a NASA type rocket or a military war headed rocket. So, I'm hoping that it's the NASA version. But, looking at the 200kg, that seems to be a bit low for one of those things launched from Cape Kennedy or wherever they launch these thing from. Bon voyage http://perendis.webs.com

OpenStudy (irishboy123):

if you didn't do so already, it's always a good idea to use a FBD